# Transferable Graph Learning for Transmission Congestion Management via Busbar Splitting

### v1

@@ Abstract
Network topology optimization (NTO) via busbar splitting can mitigate transmission grid congestion and reduce redispatch costs. However, solving this mixed-integer non-linear problem for large-scale systems in near-real-time is currently intractable with existing solvers. Machine learning (ML) approaches have emerged as a promising alternative, but they have limited generalization to unseen topologies, varying operating conditions, and different systems, which limits their practical applicability. This paper formulates NTO for congestion management problem considering linearized AC PF, and proposes a graph neural network (GNN)-accelerated approach. We develop a heterogeneous edge-aware message passing NN to predict effective busbar splitting actions as candidate NTO solutions. The proposed GNN captures local flow patterns, achieves generalization to unseen topology changes, and improves transferability across systems. Case studies show up to $10^{4}$ speed-up, delivering AC-feasible solutions within one minute and a 2.3% optimality gap on the GOC 2000-bus system. These results demonstrate a significant step toward near-real-time NTO for large-scale systems with topology and cross-system generalization.
@@
Axillary variables for power generation.
@@
Axillary variables for load demand.
@@ I Introduction
Congestion management (CM) in transmission networks is becoming increasingly challenging due to the growing integration of distributed energy resources and electrification across sectors. NTO, particularly substation reconfiguration via busbar splitting, is a promising approach to relieve congestion and reduce re-dispatch costs. However, solving this mixed-integer non-linear optimization problem for large-scale systems in near-real-time remains computationally challenging. As a result, operators often rely on expert knowledge or predefined manuals, which can lead to sub-optimal system performance [1]. Recent approaches have explored ML to identify effective topological actions in near-real-time, but most models lack scalability, remain system-specific and fail to generalize to new topologies, operating conditions, or unseen systems. This lack of generalization and transferability limits their practical use due to: 1) the need for retraining and costly data collection for each new system, 2) difficulty for operators to detect when models become outdated, and 3) performance drops under rapidly changing grid conditions. In this context, this paper proposes a GNN-based approach that exploits the locality of CM to scale to large-scale systems, generalize to topology changes, and transfer to unseen grids.
@@ II Congestion Management via Busbar Splitting
$-z_{l,i}\overline{S}_{l}\leq P_{l,i,b2}\leq z_{l,i}\overline{S}_{l}$ (2g)
@@
where (4a)-(4b) present power balance for each substation’s busbars. Active and reactive power flows are defined in (4c)-(4d), with $P^{L}_{ij,c},Q^{L}_{ij,c}$ representing active and reactive power losses. (4e) models the thermal limit of lines. (4f) represents voltage limits. We use a piecewise linearization of quadratic constraint (4e), and a linear approximation of loss terms $P^{L}_{ij,c},Q^{L}_{ij,c}$ around an initial operating point as in [21].
@@
where $\gamma\geq 1$ is a parameter, and $\Phi(.)$ is a convex function. The objective penalizes normalized line apparent flows (i.e., current magnitudes) [5]. In this work, we use use $\Phi(.)=\text{clip}(.,s^{th},\infty)$, to restrict penalization to lines whose loading exceeds a specified threshold $s^{th}$ (e.g., 80%), whereas the parameter $\gamma\geq 2$ shapes the penalty function by emphasizing highly loaded lines and down-weighting lightly loaded ones. We apply a piecewise linearization of the objective to obtain a MILP formulation.
@@ III-A GNN-accelerated NTO approach
Fig. 3 depicts the proposed GNN-NTO approach. First, given a congested grid, a proximity filter is applied to mask substations located within $k$ hops (e.g., 5 hops) of the congested lines, and with at least four line connections (3c)-(3d), denoted as $\mathcal{V}^{p}$. Notably, no input features are removed and all information is propagated locally through the GNN. The proximity filter instead limits the output nodes to substations likely to affect congestion ($\mathcal{V}^{p}$), improving learning performance. Next, the grid features are encoded and passed through the proposed GNN, which predicts the splitting status of the masked substations $\mathcal{V}^{p}$. The top-$x$ (e.g., top-3) candidates are then selected based on GNN predictions and used as potential splitting actions for the AC-NTO-CM problem, while the remaining substations are modeled as single nodes. These candidate substations are also prioritized as branching variables in the branch-and-bound MIP solver [22]. As a result, the proposed approach substantially accelerates the NTO-CM.
@@ III-B Learning Task
This paper investigates two learning objectives: (1) whether to split a node, modeled as classification (Clf), and (2) how much congestion is reduced by splitting a node, modeled as regression (Reg).
@@
In the Reg task, the goal is to predict how much congestion reduction is achieved by splitting substation $i$, expressed as $y_{i}^{Reg}=\Phi(\Delta\mathcal{O}^{CM}_{i})$, where $\Phi$ is a convex monotonic function. The Reg task provides a richer estimate of action effectiveness, allowing better interpretability and prioritization of candidate actions by operators/MIP solvers. However, it requires hyperparameter tuning of $\gamma$ in (5) and selecting $\Phi$. In this work, we apply a clipping function for $\Phi$ to limit the influence of extreme values and stabilize the learning.
@@ IV-D Transferability across Systems
Each power system has a distinct input and output distribution, posing a fundamental challenge for ZS transferability, as models trained on one system fail to generalize directly to another. However, in the TL case, retraining only the encoder and decoder with a limited number of samples effectively mitigates the distribution shift, while preserving the learned GNN embeddings. As system size increases, models encounter more diverse topologies and feature patterns, which improves TL performance, achieving an F1-score of up to 0.92 for the GOC-500 and GOC-2000 systems. Finally, the proposed MPNN achieves the best results in the CD case, highlighting its capacity to capture local PF patterns that improve cross-system generalization. This result show that, instead of generating extensive training data for large-scale grids (which is computationally expensive), one can efficiently train a universal GNN using data from decomposed, congestion-prone regions of the grid.
@@ IV-E Computational Efficiency
Table IV compares the computational efficiency of different NTO-CM approaches over three cases: DC PF with maximum 1 splitting, DC PF with maximum 3 splittings, and AC PF with maximum 1 splitting. The MIPGap in Gurobi is set to 1%, and the time limit is 10 hours. Top 5 predictions are used as candidates in GNN-NTO. The high optimality gap of the No-SW indicates the effectiveness of busbar splitting in reducing grid congestion (up to 10%). The negligible optimality gap of under 1% for the 5-hop-NTO confirms that considering the proximity filter and exploiting the locality property of the NTO-CM problem is effective, especially in larger systems where solving the full optimization becomes intractable. Moreover, the GNN-NTO approaches achieve substantial computational speed-ups, up to 4 orders of magnitude for the GOC-2000 system, while the Reg showing lower optimality gaps than the Clf due to its ability to prioritize more impactful actions with the proposed index. In the multiple switching case, the GNN-NTO optimality gap increases up to maximum 6% as the GNN is trained to predict the effectiveness of a single splitting action and may miss multiple cooperative actions. Finally, the computational advantage of GNN-NTO becomes even more pronounced for large-scale systems under AC PF, where even the proximity-filtered optimization is intractable; in such cases, GNN-NTO can still provide feasible solutions within a minute and 2.3% optimality gap. With more powerful hardware resources and efficient implementation (e.g., in lower-level languages or on GPUs), MIP process can be further reduced to a few seconds, achieving near-real-time performance.
@@ V Conclusion and Future Work
This paper presents a GNN-accelerated approach (GNN-NTO) for congestion management via busbar splitting. The GNN-NTO significantly reduces computations, enabling near-real-time solutions with operational feasibility. To this end, we develop a linear AC NTO formulation to reduce grid congestion. A novel edge-aware MPNN is proposed to predict effective busbar splitting actions, which serve as candidate solutions for the MIP problem. Two learning tasks are explored: a classification task that predicts breaker statuses, and a regression task that estimates the impact of each splitting action, providing interpretability for GNN’s predictions. A proximity filter is also used to improve learning performance by focusing on the most relevant local regions. Case studies demonstrate that the proposed MPNN effectively learns generalizable features from local PF patterns, achieving generalization to unseen topology changes and enabling cross-system transfer, where a pre-trained GNN can be efficiently transferred to new grids or a universal GNN can be trained on multiple grids. These advantages are particularly important for large-scale systems with frequent topology changes, varying grid conditions, and expensive data collection. Moreover, the GNN-NTO achieves up to 4 orders of magnitude speed-ups for the large-scale GOC 2000-bus system under AC PF, providing a feasible solution in one minute with a minor optimality gap.
Future research will focus on zero-shot transfer to unseen grids, addressing distribution shifts in ML-based optimization proxies, and extending the approach for sequential topological reconfigurations.

### v2

@@ Abstract
Network topology optimization (NTO) via busbar splitting can mitigate transmission grid congestion and reduce redispatch costs. However, solving this mixed-integer nonlinear problem for large-scale systems in near-real-time is currently intractable with existing solvers. Machine learning (ML) approaches have emerged as a promising alternative, but they have limited generalization to unseen topologies, varying operating conditions, and different systems, which limits their practical applicability. This paper formulates NTO for congestion management considering linearized AC power flow, and proposes a graph neural network (GNN)-accelerated approach. We develop a heterogeneous edge-aware message passing GNN to predict effective nodes for busbar splitting actions as candidate NTO solutions. The proposed GNN captures local flow patterns, improves generalization to unseen topology changes, and enhances transferability across systems. Case studies show up to $10^{4}\times$ speed-up, delivering AC-feasible solutions within one minute and a 2.3% optimality gap on the GOC 2000-bus system. These results demonstrate a significant step toward near-real-time NTO for large-scale systems with topology and cross-system generalization.
@@
$\mathcal{L}^{cong}$
Set of congested transmission lines.
$\mathcal{V}^{p}$
Set of substations selected by the proximity filter.
$\mathcal{V}^{cnd}$
Set of candidate substations selected by the GNN for potential busbar splitting.
@@
Auxiliary variables for power generation.
@@
Auxiliary variables for load demand.
@@ I Introduction
Congestion management (CM) in transmission networks is becoming increasingly challenging due to the growing integration of distributed energy resources and electrification across sectors. NTO, particularly substation reconfiguration via busbar splitting, is a promising approach to relieve congestion and reduce re-dispatch costs. However, solving this mixed-integer non-linear optimization problem for large-scale systems in near-real-time remains computationally challenging. As a result, operators often rely on expert knowledge or predefined manuals, which can lead to sub-optimal system performance [15]. Recent approaches have explored ML to identify effective topological actions in near-real-time, but most models lack scalability, remain system-specific and fail to generalize to new topologies, operating conditions, or unseen systems. This lack of generalization and transferability limits their practical use due to: 1) the need for retraining and costly data collection for each new system, 2) difficulty for operators to detect when models become outdated, and 3) performance drops under rapidly changing grid conditions. In this context, this paper proposes a GNN-based approach that exploits the locality of CM to scale to large-scale systems, generalize to topology changes, and transfer to unseen grids [24].
@@
The paper is organized as follows. Section II introduces the proposed NTO MILP formulation. Section III presents the GNN-based solution approach. Section IV presents the case studies, and Section V concludes the paper.
@@ II Congestion Management via Busbar Splitting
$-z_{l,i}\overline{S}_{l}\leq P_{l,i,b2}\leq z_{l,i}\overline{S}_{l}\qquad\qquad\qquad\qquad$ (2g)
@@
where (4a)-(4b) present power balance for each substation’s busbars. Active and reactive power flows are defined in (4c)-(4d), with $P^{L}_{ij,c},Q^{L}_{ij,c}$ representing active and reactive power losses. (4e) models the thermal limit of lines. (4f) represents voltage limits. We use a piecewise linearization of quadratic constraint (4e), and a linear approximation of loss terms $P^{L}_{l,i},Q^{L}_{l,i}$ around an initial operating point as in [1].
@@
where $\gamma\geq 1$ is a parameter, and $f(.)$ is a convex function. The objective penalizes normalized line apparent flows (i.e., current magnitudes) [3]. In this work, we use $f(.)=\text{clip}(.,s^{th},\infty)$, to restrict penalization to lines whose loading exceeds a specified threshold $s^{th}$ (e.g., 80%), whereas the parameter $\gamma\geq 2$ shapes the penalty function by emphasizing highly loaded lines and down-weighting lightly loaded ones. We apply a piecewise linearization of the objective to obtain a MILP formulation.
@@ III-A GNN-accelerated NTO approach
Fig. 3 and Algorithm 1 present the workflow of the proposed approach, consisting of three stages: proximity filter, GNN-based candidate prediction, and accelerated NTO.
III-A1 Proximity Filter
Given a congested grid state, the first step identifies substations likely to influence congestion. Congested lines $\mathcal{L}^{cong}$ are detected based on a threshold $s^{th}_{l}$ (Line 1). A proximity filter then selects substations within $k$ hops (e.g., 5 hops) of the congested lines and with at least four connected lines, as defined in (3c)–(3d), forming the proximity node set $\mathcal{V}^{p}$ (see Fig. 2). The filter does not remove input features; instead, it restricts the GNN output space to substations more likely to affect congestion, improving learning efficiency. During training, backpropagation is also applied only to these nodes.
III-A2 GNN-Based Candidate Prediction
The grid state is encoded as a graph and processed by the proposed GNN, which predicts a splitting score $\hat{y}_{i}$ for each substation $i\in\mathcal{V}^{p}$ (Line 1). These scores indicate how promising splitting a substation is for congestion mitigation. Substations are then ranked based on $\hat{y}_{i}$, and the top-$x$ (e.g., top-3) are selected as candidate nodes $\mathcal{V}^{cnd}$ (Line 1).
III-A3 Accelerated NTO
The predicted candidate substations are incorporated into the NTO optimization. Substations outside the candidate set are fixed as unsplit nodes ($z_{i}=1\text{ }i\notin\mathcal{V}^{p}$), significantly reducing the number of binary variables (Line 1). Moreover, candidate variables $z_{i}\text{ }i\in\mathcal{V}^{p}$ are prioritized as branching variables in the branch-and-bound (B&B) process of the MIP solver [10] (Line 1). In other words, these variables are assigned higher branching priority compared to other binary variables (e.g., $z_{l,i},z_{g},z_{d}$), guiding the solver to explore promising topology decisions earlier in the search process. The reduced problem, denoted NTO, is then solved using a commercial solver (Line 1).
By focusing the search on high-quality candidates predicted by the GNN, the proposed approach significantly accelerates the NTO-CM solution while maintaining high solution quality.
1
Input: $\hat{P}_{d},\hat{Q}_{d}$ $d\in\mathcal{D}$, $\hat{P}_{g},\hat{Q}_{g}$ $g\in\mathcal{G}$, $P_{l},Q_{l}$ $l\in\mathcal{L}$
2
3 Detect congested lines $\mathcal{L}^{cong}=\{l\in\mathcal{L}\mid S_{l}\geq s^{th}\}$;
4
Proximity Filter
5
6for $l\in\mathcal{L}^{cong}$ do
7 for $i\in\mathcal{V}$ do
8    if $dist(l,i)\leq k$ and $Deg(v)\geq 4$ then
9         add $i$ to $\mathcal{V}^{p}$;
10
11
12
GNN-Based Candidate Prediction
13 Forward pass of GNN $\hat{y}_{i}=\phi^{GNN}(x,e)\quad i\in\mathcal{V}^{p}$;
14
15Select the top-$x$ substations $i$ from $\mathcal{V}^{p}$ ranked by $\hat{y}_{i}$, and add them to $\mathcal{V}^{cnd}$;
16
Accelerated NTO;
17
18Fix non-candidate substations $z_{i}=1\quad i\notin\mathcal{V}^{cnd}$ ;
19
20Prioritize candidate variables $z_{i}\quad i\in\mathcal{V}^{cnd}$ in B&B;
21
22Solve reduced NTO problem (5);
23 return optimal topology;
Algorithm 1 GNN-NTO Approach
@@ III-B Learning Task
This paper investigates two learning objectives: (1) whether to split a node, modeled as classification (Clf), and (2) how much congestion is reduced by splitting a node, modeled as regression (Reg). In both settings, the effect of splitting substation $i$ is evaluated independently, assuming that all other substations remain unsplit.
@@
In the Reg task, the goal is to predict how much congestion reduction is achieved by splitting substation $i$, expressed as $y_{i}^{Reg}=\Phi(\Delta\mathcal{O}^{CM}_{i})$, where $\Phi$ is a convex monotonic function. The Reg task provides a richer estimate of action effectiveness, allowing better interpretability and prioritization of candidate actions by operators/MIP solvers. However, it requires hyperparameter tuning of $\gamma$ in (5) and selecting $\Phi$. In this work, we use a clipping function $\Phi(.)=\text{clip}(.,\Delta\mathcal{O}^{low}_{i},\Delta\mathcal{O}^{high}_{i})$ to limit the influence of extreme values and stabilize the learning.
@@ III-C2 Heterogeneous GNN
III-D Training Data Generation
Algorithm 2 summarizes the training data generation process. First, random demands, planned NkN\!-\!k line outages, and generation costs are sampled to emulate diverse operating conditions. The LP-AC-OPF is then solved to obtain the generation dispatch and initial power flows. To improve learning performance, only congested grid states are retained for training [18].
Next, the proximity filter is applied to determine the proximity node set $\mathcal{V}^{p}$. For each $i\in\mathcal{V}^{p}$, (6) is solved to evaluate the effectiveness of splitting substation $i$ for congestion management. Since (6) only involves elements of substation $i$ and assumes other substations as single electric nodes, these problems can be solved very efficiently and in parallel for all $i\in\mathcal{V}^{p}$. The label is then assigned based on the prediction task.
1
2for data sample $s\in\Omega^{s}$ do
3
4  Sample inputs: $\hat{P}_{d,s},\hat{Q}_{d,s}$ $d\in\mathcal{D}$, outages $l\in\Omega^{o}_{s}$, generation costs $c_{g}$ $g\in\mathcal{G}_{s}$;
5
6  Solve LP-AC-OPF to obtain $\hat{P}_{g},\hat{Q}_{g}$ $g\in\mathcal{G}$;
7
8  Compute base congestion cost $O^{CM}_{0}$;
9
10 if no congestion then
11      continue to next sample;
12
13
14  Apply proximity filter to obtain $\mathcal{V}^{p}$;
15
16 for $i\in\mathcal{V}^{p}$ do
17      Solve (6) for $z_{i},z_{e\in E_{i}},\mathcal{O}^{CM}_{i}$;
18
19     Compute congestion reduction $\Delta\mathcal{O}^{CM}_{i}$;
20
21     Assign label $y_{i}$ according to Section III-B;
22
23
Algorithm 2 Training Data Generation
@@ IV-D Transferability across Systems
Each power system has a distinct input and output distribution, posing a fundamental challenge for ZS transferability, as models trained on one system fail to generalize directly to another. However, in the TL case, retraining only the encoder and decoder with a limited number of samples effectively mitigates the distribution shift, while preserving the learned GNN embeddings. As system size increases, models encounter more diverse topologies and feature patterns, which improves TL performance, achieving an F1-score of up to 0.92 for the GOC-500 and GOC-2000 systems. Finally, the proposed MPNN achieves the best results in the CD case, highlighting its capacity to capture local PF patterns that improve cross-system generalization. This result shows that, instead of generating extensive training data for large-scale grids (which is computationally expensive), one can efficiently train a universal GNN using data from decomposed, congestion-prone regions of the grid.
@@ IV-E Computational Efficiency
Table IV compares the computational efficiency of different NTO-CM approaches over three cases: DC PF with maximum 1 splitting, DC PF with maximum 3 splittings, and AC PF with maximum 1 splitting. The MIPGap in Gurobi is set to 1%, and the time limit is 10 hours. Top 5 predictions are used as candidates in GNN-NTO. The high optimality gap of the No-SW indicates the effectiveness of busbar splitting in reducing grid congestion (up to 10%). The negligible optimality gap of under 1% for the 5-hop-NTO confirms that considering the proximity filter and exploiting the locality property of the NTO-CM problem is effective, especially in larger systems where solving the full optimization becomes intractable. The GNN-NTO approaches achieve substantial computational speed-ups, up to 4 orders of magnitude for the GOC-2000 system. In this context, the Reg model achieves slightly lower optimality gaps than the Clf model, as it can better prioritize impactful splitting actions using the proposed congestion reduction index. However, this improved ranking performance comes with increased sensitivity to hyperparameters and training design. The practical differences between the Clf and Reg tasks are discussed in the next section. In the multiple switching case, the GNN-NTO optimality gap increases up to maximum 6% as the GNN is trained to predict the effectiveness of a single splitting action and may miss multiple cooperative actions. Finally, the computational advantage of GNN-NTO becomes even more pronounced for large-scale systems under AC PF, where even the proximity-filtered optimization is intractable; in such cases, GNN-NTO can still provide feasible solutions within a minute and 2.3% optimality gap. With more powerful hardware resources and efficient implementation (e.g., in lower-level languages or on GPUs), MIP process can be further reduced to a few seconds, achieving near-real-time performance.
@@
IV-F Discussion
The presented case studies demonstrate promising results for the proposed GNN-NTO approach for congestion management. Table I shows that the MPNNs improve prediction accuracy compared to other architectures, while the heterogeneous GNN model does not provide additional benefit for the node candidate prediction task in this work. Table II shows that generalization to topology changes improves in larger systems, although zero-shot generalization to unseen topologies remains limited. Table III shows that the local formulation of the learning task together with the MPNN design improves transferability across systems with only minimal additional training or by training a universal model, while zero-shot generalization to entirely unseen systems remains a challenge. Finally, Table IV demonstrates the substantial computational speed-up achieved by the GNN-NTO approach while maintaining relatively small optimality gaps. However, some limitations of our approach can be noted.
Regarding the learning tasks, the Reg provides a more informative signal for ranking candidate substations. However, it is also more sensitive to the choice of hyperparameters $\gamma$ and $\Phi$ that define the congestion reduction index. Large values of $\gamma$ emphasize highly congested lines but may increase sensitivity to small variations in the grid state. In addition, the regression index can take negative values when a split worsens congestion, which can bias the learning objective if not properly controlled. Therefore, clipping in $\Phi$ is used for stable training. In this work, the regression formulation is explored as a proof-of-concept for predicting congestion reduction magnitudes to improve candidate selection. Further work is needed to refine this approach.
The GNN predicts the effectiveness of a busbar splitting action while assuming that no other substations are split. Consequently, when multiple splitting actions are required to mitigate congestion, the current approach identifies them sequentially, which does not guarantee finding the optimal combination. In addition, although the GNN-NTO approach provides significant computational speed-ups during online use, generating labeled training data remains computationally expensive for large-scale systems. Improving zero-shot transferability can address this limitation by enabling models trained on smaller systems to be applied directly to larger systems.
@@ V Conclusion and Future Work
This paper presents a GNN-accelerated approach (GNN-NTO) for congestion management via busbar splitting. The proposed approach significantly reduces computations, enabling near-real-time solutions with operational feasibility. To this end, we develop a linearized AC NTO formulation to reduce grid congestion. A novel edge-aware MPNN is proposed to predict effective nodes for busbar splitting actions, which serve as candidate solutions for the MIP problem. A proximity filter further improves learning performance by focusing the prediction task on the most relevant local regions. Two learning tasks are explored: a classification task that predicts whether splitting a substation alleviates congestion, and a regression task that estimates the congestion reduction associated with each splitting action. While the regression formulation can improve the prioritization of candidate nodes and provide interpretability, it is also more sensitive to hyperparameter choices and requires further investigation. Case studies demonstrate that the proposed MPNN learns generalizable features from local PF patterns. In particular, the results show that generalization to topology changes improves in larger systems, and that transfer learning across systems is effective with only minimal additional training, allowing a pre-trained GNN to be efficiently adapted to new grids or enabling a universal model to be trained across multiple systems. These properties are particularly valuable for large-scale systems with frequent topology changes, varying grid conditions, and expensive data collection. However, zero-shot generalization to entirely unseen systems remains challenging due to distributional shifts in the input and output spaces. Moreover, the GNN-NTO achieves up to 4 orders of magnitude speed-ups for the large-scale GOC 2000-bus system under AC PF, providing a feasible solution in one minute with a minor optimality gap.
Future research will focus on zero-shot generalization to unseen grids, addressing distributional shifts in ML-based optimization proxies, and extending the approach for multiple busbar splitting actions.
Acknowledgment
This work is supported by the TU Delft AI Labs & Talent Programme. The authors are thankful to Olayiwola Arowolo from Delft University of Technology who provided valuable discussions.